OpenStudy (anonymous):

(3.3 x 10^6) divided by (3 X 10^8)

OpenStudy (anonymous):

just calculate and divide

OpenStudy (anonymous):

do u have a ti calculator?

OpenStudy (anonymous):

no i don't have a ti calculator

OpenStudy (anonymous):

.011

OpenStudy (anonymous):

u need to purchase a ti calculator it will help

OpenStudy (anonymous):

You can always treat the two parts of scientific notation separately when multiplying or dividing. So, 3.3/3 = 1.1 and 10^6 / 10^8 = 10^-2. Thus, (3.3 x 10^6)/(3 x 10^8) = 1.1 x 10^-2.